OpenStudy (anonymous):

Please Someone solve it integration of(1-v)/(1+V`2) v`2=V Square denominator v is square

OpenStudy (anonymous):

i think you have to first separate the integral: 1/(1+v^2) + v/(1+v^2), so the integral of the first term is ArcTan (v) and for the integral of the second term i recommend you to do a change of variable: y= 1+v^2 so dy= 2v and v= dy/2 and then you substitute: v/(1+v^2) = (1/2)(dy/y) and the integral is (1/2) (In y) finally you plug in the initial variables: (1/2)(In [1+v^2]) so the total integral is: ArcTan (y) + (1/2)(In [1+v^2])